# DirectMultiStep: Direct Route Generation for Multi-Step Retrosynthesis

### v1

@@ 3.1 Test accuracy
Figure 4: Separation of a 10-step route from set-n1 into two 5-step routes. (a) Correct prediction for the first half of the 10-step route with starting material information. (b) Correct prediction for the second half of the 10-step route with starting material information.
The requirement to provide the starting material structure as input to the model is a limitation, although not as significant as it may seem. Corey’s retrosynthetic framework allows finding at least one route to starting materials by performing one functional group transformation at a time, usually requiring protection of other functional groups. However, such a route will be too long to be efficient in terms of yield (overall yield is the product of yields of individual steps). The art of organic synthesis lies in finding ways to perform several transformations simultaneously using a minimal number of protecting groups. We believe our model provides value as an assistant to this task: finding better ways to get to the target compound from a starting material. The requirement to specify the number of steps as an input becomes a useful option in this context. Moreover, if the target compound is chiral (i.e., it has a mirror image which cannot be superimposed), even during manual retrosynthetic planning, chemists check if chiral centers can be obtained from chiral starting materials (such as naturally occurring amino acids) because developing reactions that create chiral centers in the desired conformation remains an active research area (and subject of 2001 and 2021 Nobel Prizes in Chemistry). An example of such a route is shown in Fig. 6.
@@ 3.2 Separation of Long Routes into Shorter Routes
As seen in Fig. 3, our model does not find the single 10-step route in the n1 set. However, this route can be correctly reproduced with rank 1 if split into two 5-step routes and the intermediate compound is provided as the starting material for the first 5 reactions (Fig. 4) and a target compound for the last 5 reactions (Fig. 4). While ideally the necessity to provide an intermediate compound should be avoided, this resembles the bidirectional search technique commonly employed in manual retrosynthetic analysis [44].
3.3 Retrosynthetic planning of pharmaceutical compounds
We demonstrate the generalizability of our model by testing on three FDA-approved drugs: Vonoprazan, Mitapivat, and Daridorexant, which were used for evaluation by Xiong, et al. [45]. These drugs and their intermediates are absent from the training set. Vonoprazan, a potassium competitive acid blocker for Helicobacter pylori infections, was initially proposed with a 2-step route (Fig. 5[46]. However, one of its starting materials is unstable, and the final reaction leads to significant byproducts. Subsequently, an improved 4-step route (Fig. 5) was introduced [47]. Mitapivat, a pyruvate kinase activator for treating hemolytic anemia, has been associated with both a 5-step route (Fig. 5) and a 3-step route (Fig. 5[48, 49]. Daridorexant, an orexin receptor antagonist for adult insomnia, is linked to a 4-step route (Fig. 6[50].
@@ 3.3 Retrosynthetic planning of pharmaceutical compounds
Figure 5: Literature routes for Vonoprazan and Mitapivat correctly reproduced by our model. Target compounds are in green, starting materials that are given as inputs are colored in orange. Ranks denote the rank of this route when the specified starting material is provided. (a) First literature route for Vonoprazan. The model predicts correctly no matter which starting material is given. (b) Second literature route for Vonoprazan. The model predicts the route correctly only when an immediate precursor to Vonoprazan is given as the target compound. (c) First literature route for Mitapivat. (d) Second literature route for Mitapivat.
@@
Our model correctly reproduces Vonoprazan’s first literature route (Fig. 5) regardless of the starting material provided, with high ranks. However, the model struggles to find the first transformation (counting from the target compound) from the second route. In the first route, the first step is a reductive amination, while in the second route, it’s a reduction of the carbonyl bond in the amide group. Reductive amination is much more common for creating C-N bonds, so the training set is heavily biased toward it. As a result, all routes predicted by our model start as in Fig. 5. However, if we provide the precursor to Vonoprazan in the second route as the target compound, the model correctly reproduces the full route (Fig. 5). This highlights an important limitation of our model: it may not perform as well on routes involving transformations that are inadequately represented in the training dataset. We envision that as larger multi-step datasets become available, this problem will naturally disappear.
@@
Figure 6: Daridorexant literature multi-step routes and predictions from the model with starting material information. The compounds that are missing from the prediction are in red.
@@
Our model correctly reproduces both routes for Mitapivat (Fig. 5 and Fig. 5). In contrast, the route for Daridorexant (Fig. 6) is only partially predicted: our model does not split the right child of the root node (Daridorexant itself) into the two compounds highlighted in red. All other intermediates and starting materials are predicted correctly. The reason for this behavior is not well understood, but it can be rationalized by looking at the distribution of the number of leaves at root nodes: 73% of root nodes (target compounds) have at least 1 leaf (see SI Fig. 7). Such dependence on the distributions of the training set comprises an important limitation of the multi-step first approach. This issue would not occur with methods based on single-step prediction because the tree search framework would not halt at compounds outside the stock compound set before reaching the maximum iterations. Ensuring that multi-step first models reproduce that aspect of SSR methods constitutes a direction for future research.
@@ 4 Conclusion
In this work, we introduced DirectMultiStep, a transformer-based model that directly generates multi-step synthetic routes for a given target compound. By predicting the entire route as a single string, our approach bypasses the need for iterative single-step predictions and exponential search space traversal. The model accommodates specific conditions such as the desired number of steps and starting materials, enabling efficient route planning.
DirectMultiStep outperforms state-of-the-art methods on the PaRoutes dataset, achieving a 2.2x improvement in Top-1 accuracy on the n1 test set and a 3.3x improvement on the n5 test set. The model also successfully predicts routes for FDA-approved drugs not included in the training data, showcasing its generalization capabilities. However, the model’s performance may be impacted by the suboptimal diversity of the current training set, particularly for less common reaction types.
Despite these limitations, the DirectMultiStep approach presents a promising direction towards fully automated retrosynthetic planning. As larger and more diverse multi-step datasets become available, the model’s performance is expected to improve further. Future research should focus on addressing the limitations identified, such as the need to split longer routes into sub-routes and the dependence on the distributions of the training set. Overall, this work demonstrates the potential of transformer-based models for efficient and scalable computer-aided synthesis planning.
5 Broad Impact
Chemical synthesis of target compounds of interest forms the routine for millions of organic chemists around the world, including graduate students in PhD programs and synthetic chemists in pharmaceutical companies. The ability to find shorter (and thus more efficient) routes to any desired compound is an enticing challenge with the potential of saving on labor costs and chemical waste. As a result, multi-step prediction methods have the potential to improve the lives of millions of people, and potentially contribute to faster advances in the development of new drugs and materials. We hope that our work constitutes an important advancement in that direction.
Quite unfortunately, however, there is no fundamental chemical difference between how safe and useful compounds (such as drugs) and malign compounds (such as chemical weapons) are synthesized. Both require application of the same set of rules of organic synthesis, and so the ability to make drugs faster and more efficiently inherently brings along the ability to make poisons faster and more efficiently. One cannot impose restrictions on the tokens (chemical elements) learned by the model, as even some combinations of nitrogen and oxygen (which are widespread in proteins and DNA) may be cytotoxic. Therefore, unless the humanity at large decides that the benefits of automated retrosynthesis do not outweigh its potential harms, the most we can do is communicate the dangers clearly and perhaps impose restrictions that would still require a trained (which often accompanies training in ethics) chemist to use our model. Besides, all meta-data of reaction conditions and catalysts are removed from the predictions of our model, which would present a deliberate barrier against malicious uses.
6 Code and Data Availability
@@
The reaction route dataset is from the 2.0 version of PaRoutes in their GitHub repository [39], available under the Apache License 2.0. Code to process dataset, implementation of model architecture, code for training, generation, and evaluation are available under MIT License at https://github.com/batistagroup/DirectMultiStep.
7 Acknowledgments

### v2

@@ 3.1.2 MoE Transformers
Figure 4: Separation of a 10-step route from set-n1 into two 5-step routes. (a) Correct prediction from DMS-Flex (Duo) for the first half of the 10-step route with starting material information (red). (b) Correct prediction from DMS-Flex (Duo) for the second half of the 10-step route with starting material information.
3.1.3 Use Cases for Different DMS Model Variants
The requirement to provide the starting material structure as input to the model might, in theory, seem as a limitation, limiting the relevance of DMS-Flash/Flex models. However, in practice, a chemist often can know the structure of a desired starting material because Corey’s retrosynthetic framework allows finding at least one route to starting materials by performing one functional group transformation at a time. Such deterministically found routes will often be too long to be efficient in terms of yield (overall yield is the product of yields of individual steps). The art of organic synthesis lies in finding ways to perform several transformations simultaneously using a minimal number of protecting groups. As such, we believe our SM-based models provide value as an assistant to this task: finding better ways to get to the target compound from a starting material. The requirement to specify the number of steps as an input becomes a useful option in this context. Moreover, if the target compound is chiral (i.e., it has a mirror image which cannot be superimposed), even during manual retrosynthetic planning, chemists check if chiral centers can be obtained from chiral starting materials (such as naturally occurring amino acids) because developing reactions that create chiral centers in the desired conformation remains an active research area (and subject of 2001 and 2021 Nobel Prizes in Chemistry). An example of such a route is shown in Fig. 7.
In addition, one can use more computationally expensive DMS-Wide/Deep models to discover potential starting materials, and then use those structures as inputs to DMS-Flash/Flex to generate suggestions for each starting material of interest.
To lift the necessity to specify desired route length, a user also has a choice between running Flex/Wide models several times with different desired route lengths or simply running Explorer-XL once. Of course, ideally one would prefer a single model that predicts the whole route with just from the structure of the target compound, without number of steps or starting material provided as an extra input. We believe better performance could be attainable in the future once bigger datasets become available. Finding ways to improve performance even with current dataset constitutes an important area of further research.
@@ 3.2 Separation of Long Routes into Shorter Routes
As seen in Fig. 3, DMS-Flex (Duo) does not find the single 10-step route in the n1 set. However, this route can be correctly reproduced with rank 1 if split into two 5-step routes and the intermediate compound is provided as the starting material for the first 5 reactions (Fig. 4) and a target compound for the last 5 reactions (Fig. 4). While ideally the necessity to provide an intermediate compound should be avoided, this resembles the bidirectional search technique commonly employed in manual retrosynthetic analysis [49]. In contrast, models trained without SM, such as DMS-Deep and DMS-Wide, can replicate the second 5-step sequence in Fig. 4 but not the first (Fig. 4).
@@
Figure 5: Literature routes for Vonoprazan and Mitapivat correctly reproduced by DMS-Flex (Duo). Target compounds are in blue, starting materials that are given as inputs are colored in red. Ranks denote the rank of this route when the specified starting material is provided. (a) First literature route for Vonoprazan. The model predicts correctly no matter which starting material is given. (b) Second literature route for Vonoprazan. The model predicts the route correctly only when an immediate precursor to Vonoprazan is given as the target compound. (c) First literature route for Mitapivat. (d) Second literature route for Mitapivat.
3.3 Retrosynthetic planning of pharmaceutical compounds
For another generalizability assessment of our model, we test it on three FDA-approved drugs: Vonoprazan, Mitapivat, and Daridorexant, which were used for evaluation by Xiong, et al. [50]. These drugs and their intermediates are absent from the training set. Vonoprazan, a potassium competitive acid blocker for Helicobacter pylori infections, was initially proposed with a 2-step route (Fig. 5[51]. However, one of its starting materials is unstable, and the final reaction leads to significant byproducts. Subsequently, an improved 4-step route (Fig. 5) was introduced [52]. Mitapivat, a pyruvate kinase activator for treating hemolytic anemia, has been associated with both a 5-step route (Fig. 5) and a 3-step route (Fig. 5[53, 54]. Daridorexant, an orexin receptor antagonist for adult insomnia, is linked to a 4-step route (Fig. 7[55].
@@ 3.3 Retrosynthetic planning of pharmaceutical compounds
Both vanilla transformers (DMS-Flash) and MoE models (DMS-Flex Duo) correctly reproduce Vonoprazan’s first literature route (Fig. 5) regardless of the starting material provided, with high ranks. However, DMS-Flash struggles to find the first transformation (counting from the target compound) in the second route. This inability could be attributed to the fact that the reduction of amide carbonyl (the step leading to Vonoprazan in second route) is less frequently represented in the PaRoutes dataset than the much more common reductive amination (the step leading to Vonoprazan in first route). As a result, most routes predicted by DMS-Flash start with a reductive amination as in Fig. 5. However, if we provide the precursor to Vonoprazan in the second route as the target compound, the model correctly reproduces the remaining route (Fig. 5). This underscores that the performance of our models relies upon sufficient representation of transformations in the training set. This limitation, however, should naturally disappear once larger multistep datasets become available.
Inability of DMS-Flash to suggest reduction of amide carbonyl in Fig. 5 was the motivation for the creation of mixture-of-experts (MoE) models. We conjectured that different experts might specialize in different classes of transformations, and even if one expert dominates the generations, one could still manually force the output from a different expert. Fortunately, no such manual interventions were needed as DMS-Flex (Duo) predicted the route almost identical to the second transformation as the Top-1 candidate (Fig. 6). The only difference from the correct route is the order of transformations: in the correct route the tosyl group is introduced after amidation with methylamine, not before.
Figure 6: A variation of second Vonoprazan route predicted by DMS-Flex (Duo). The correct route is shown in Fig. 5 and differs by the order of the last steps: tosyl group is introduced after amidation, not before.
@@
Figure 7: Daridorexant literature multistep routes and predictions from DMS-Flex (Duo). The compounds that are missing from the prediction are in dotted boxes.
@@
DMS-Flex (Duo) correctly reproduces both routes for Mitapivat (Fig. 5 and Fig. 5). In contrast, the route for Daridorexant (Fig. 7) is only partially predicted: our model does not split the right child of the root node into the two compounds in the dotted boxes. All other intermediates and starting materials are predicted correctly. The reason for this behavior is not well understood, but it can be rationalized by looking at the distribution of the number of leaves at root nodes: 73% of root nodes (target compounds) have at least 1 leaf (see Fig. S3). This dependence on the distributions of the training set comprises an important limitation of our models; however, it should also be resolved once bigger datasets become available. Admittedly, this issue would not occur with current methods based on single-step prediction because the tree search framework would not halt at compounds outside the stock compound set before reaching the maximum iterations. Ensuring that multistep-first models reproduce that aspect of SSR methods constitutes a direction for future research.
This example, however, does not imply that the model struggles to predict routes with convergent branches. If one defines a convergent route as the one that has at least one node (compound) with at least two children that are not leaves, and repeats the analysis, no significant differences in Top-K accuracies (Fig. S5-S6) or in the number of routes found (Fig. S7-S8) is observed despite significant variation in the number of routes that are convergent among training routes, n1, and n5 (Fig. S3).
@@ 4 Conclusion
In this work, we introduce the first-of-its-kind inherently multistep approach for multistep retrosynthesis prediction. By predicting the entire route as a single string, our approach bypasses the need for iterative single-step predictions and exponential search space traversal. We present a family of transformer-based DirectMultiStep (DMS) models which can predict correct routes just from the structure of the target compound, but can also accommodate specific conditions such as the desired number of steps and starting materials, enabling efficient route planning.
Our flagman DMS-Flex (Duo) model, incorporating mixture-of-experts approach, outperforms state-of-the-art methods on the PaRoutes dataset, achieving a 2.5x improvement in Top-1 accuracy on the n1 test set and a 3.9x improvement on the n5 test set. We demonstrate the quality of predicted routes by testing our models on several FDA-approved drugs. Those predictions also highlight limitations caused by the suboptimal diversity of the current training set, particularly for less common reaction types. Those limitations are partially mitigated by employing mixture-of-experts techniques, making them a promising candidate for further research in retrosynthesis prediction and applications of generative models to chemistry in general.
We believe our work demonstrates the superiority of multistep-first approach to full route prediction and inspires further research into methods that reduce the exponential complexity of iterative application of single-step based methods. We are grateful to the authors of PaRoutes for the development and release of the first multistep prediction benchmark, and we hope more benchmarks based on experimentally verified routes become available. Future research may focus on improvement of performance on longer routes (e.g. eliminating the necessity to split longer routes) and (or) incorporation of additional constraints (such as cost) on the route prediction. Overall, this work demonstrates the potential of transformer-based models for efficient and scalable computer-aided synthesis planning. We release the code used for training and route prediction as a public GitHub Repo under MIT License. Our models are also available with a Web GUI at models.batistalab.com.
5 More Computational Details
5.1 Data Curation and Preparation
A second version of the PaRoutes [44] dataset containing 450k routes is stripped of all metadata and stored as recursive dictionaries representing multistep routes. The two evaluation sets, n1 and n5, are processed similarly. A training partition is created by removing all repetitions (some routes were represented multiple times with different metadata) and permutations (swapping left and right subtrees) of routes in n1 and n5 from the full dataset, resulting in 163k routes. The training dataset is augmented by adding 2 permutations for each route. To train the DMS model which takes starting material as an input, we find all starting materials (leaves) for each tree and store a combination of the target compound with each starting material as a separate entry. As a result, DMS with SM is trained on 1 340 243 inputs, and the DMS without SM on 432 684 inputs. All SMILES strings are tokenized by treating each character as a token, and the string representation of the multistep route is tokenized similarly, treating delimiters of the tree (‘smiles’, ‘children’, ,, [, ], {, and }) as separate tokens. The final vocabulary size is 52 (including start, end, and padding tokens), the largest multistep route has 1074 tokens, while the largest target compound and starting material have 145 and 135 tokens, respectively.
5.2 Model Architecture
We present a family of DirectMultiStep (DMS) models with varying parameter counts and architectural details. DMS-Flash and DMS-Deep the classical transformer encoder-decoder architecture [56], whereas DMS-Flex and DMS-Wide incorporate sparsely-gated mixture-of-experts layers [46, 47, 48]. Major parameters choices are given in model card (Table  1). All models use GeLU activations [57].
Intermediate experiments with evaluations on subsets of n1 and n5 with 50 and 500 routes reveal no benefit in increasing the number of hidden dimensions to 512 (Tables S4-S5), which could be rationalized by noticing that 256 is already almost 5x the vocabulary size. Although a feed-forward multiplier of four is more common in NLP models, we find no significant improvement in performance as compared to the feed-forward multiplier of three (Tables S6-S7). In addition, the number of encoder layers does not have to be the same as the number of decoder layers, and models with 6 encoder and 12 decoder layers (trained with SM) performed almost as good as the models with 12 encoder and decoder layers. For MoE models, we find no significant improvement upon increase of number of experts to four, and models trained with two active experts outperform those trained with one.
5.3 Training
During training, a 10% dropout rate is used in embedding, attention, and feed-forward layers. The AdamW optimizer [58, 2019adamw] is used with a learning rate scheduler that warms up from 0 to $3.5\times 10^{-4}$ for DMS-Flex and DMS-Wide, $3\times 10^{-4}$ for DMS-Flash, and $1\times 10^{-4}$ for DMS-Deep during the first 10% of training steps and undergoes cosine decay to $10\%$ of the maximal learning rate over 20 epochs. The optimal initial learning rate for each model was chosen by starting from 5e-4 and monitoring for gradient explosion (in which case the learning rate was decreased by 0.5e-4). Gradients are clipped to 1.0. In addition, each token in the input to the encoders has a 5% chance of being masked. During hyperparameter optimization, the model was validated on a 5% split from the training routes (which exclude n1 and n5 routes).
6 Data and Software Availability
@@
The reaction route dataset is from the 2.0 version of PaRoutes in their GitHub repository [44], available under the Apache License 2.0. Code to process dataset, implementation of model architecture, code for training, generation, and evaluation are available under MIT License at https://github.com/batistagroup/DirectMultiStep. A web interface is also available at models.batistalab.com.
Author Contributions
YS, HL, and AM contributed equally to this work. YS conceived the idea. YS and HL conducted preliminary model training and testing. AM introduced the mixture of experts (MoE) approach and carried out additional model training and testing. VSB provided guidance for the research. All authors participated in writing and reviewing the manuscript.
Conflict of Interest
The authors declare no conflict of interest.
7 Acknowledgement
@@
Supporting Information
Figure S1: Lengths of valid routes predicted by DMS-Explorer for targets in n1 and n5 compared to the length of the correct (experimental) route. Left: distribution of differences in lengths for all predictions for all targets. Right: distribution of average (among all predictions for a single target) differences.
Figure S2: Lengths of valid routes predicted by DMS-Flex (Duo) for targets in n1 and n5 compared to the length of the correct (experimental) route. Left: distribution of differences in lengths for all predictions for all targets. Right: distribution of average (among all predictions for a single target) differences.
Figure S3: Distribution of the relative frequencies for the number of leaves (nodes with no children) at root node.
Table S1: Number of predicted routes in n1 that pass post-processing filters.
Model Valida Correct SMb,f SMs In Stockc Correct and In Stockd,f
Flash 9998 9942 8814 8725
Flashe 9994 7660
Explorer 9997 9923 8681 8555
Explorere 9995 7418
Explorer-XLe 9998 8008
Flex (Mono) 9941 9545 8338 7948
Flex (Duo) 9995 9866 8828 8649
Flex (Duo)e 9994 7692
Deepe 9998 7957
Widee 9996 8089
aAt least one route with all valid SMILES (including target, intermediates, and SMs).
bAt least one route with the correct conditioned SMs (and with all valid SMILES).
cAt least one route with all SMs in the n1 stock set (and with all valid SMILES).
dAt least one route with the correct conditioned SMs and with all SMs in the n1 stock set (and with all valid SMILES).
eSMs not provided during inference.
fBlank indicates not relevant (models without SM conditions).
Table S2: Number of predicted routes in n5 that pass post-processing filters.
Model Valida Correct SMb,f SMs In Stockc Correct and In Stockd,f
Flash 9991 9888 8646 8487
Flashe 9991 7576
Explorer 9997 9903 8439 8281
Explorere 9998 7227
Explorer-XLe 10000 7904
Flex (Mono) 9932 9374 8037 7528
Flex (Duo) 9993 9808 8627 8382
Flex (Duo)e 9990 7605
Deepe 9997 7830
Widee 9997 7950
aAt least one route with all valid SMILES (including target, intermediates, and SMs).
bAt least one route with the correct conditioned SMs (and with all valid SMILES).
cAt least one route with all SMs in the n5 stock set (and with all valid SMILES).
dAt least one route with the correct conditioned SMs and with all SMs in the n5 stock set (and with all valid SMILES).
eSMs not provided during inference.
fBlank indicates not relevant (models without SM conditions).
Table S3: Specifications of architecture for different model variants.
DMS Model Encoder Decoder Hidden FF Total Active Experts
Variant Layers Layers Dimension Multipliera Experts Training Generation
Van-19 6 6 256 3 1 1 1
Van-60 8 8 512 4 1 1 1
MoE-1/3 6 6 256 3 3 1 1
a an integer by which the hidden dimension is multiplied in feed-forward layers after attention.
Figure S4: Fraction of routes (grouped by route length) that are convergent in training set, n1, and n5. A route is considered convergent if any node has at least two children that are not leaves. Overall, convergent routes constitute 9.4% of the training routes, 5.6% of n1, and 10.9% of n5.
Figure S5: Top-K accuracy on the n1 test set for different models for (a) all routes in n1, (b) convergent routes in n1, and (c) non-convergent routes in n1. A route is considered convergent if any node has at least two children that are not leaves.
Figure S6: Top-K accuracy on the n5 test set for different models for (a) all routes in n5, (b) convergent routes in n5, and (c) non-convergent routes in n5. A route is considered convergent if any node has at least two children that are not leaves.
Figure S7: Distribution of the average number of routes for targets in n1 test set (a-c) after checking for the validity of smiles, (d-f) after canonicalization of SMILES and removing repetitions (considering permutations of children), and (g-i) after filtering for routes that have all reactants in stock. Distribution is shown separately for all targets (a, d, g), for targets for which an experimental route is convergent (b, e, h) and for targets for which an experimenal route is non-convergent (c, f, i). A route is considered convergent if any node has at least two children that are not leaves.
Figure S8: Distribution of the average number of routes for targets in n5 test set (a-c) after checking for the validity of smiles, (d-f) after canonicalization of SMILES and removing repetitions (considering permutations of children), and (g-i) after filtering for routes that have all reactants in stock. Distribution is shown separately for all targets (a, d, g), for targets for which an experimental route is convergent (b, e, h) and for targets for which an experimenal route is non-convergent (c, f, i). A route is considered convergent if any node has at least two children that are not leaves.
Table S4: Top-$K$ accuracy on route test set-n1 (10 000 routes).
DMS SM SM
Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
Van-19 19M 0.44 0.53 0.57 0.59 0.60 0.63
MoE-1/3 38M 0.40 0.47 0.50 0.51 0.52 0.53
Van-60 60M 0.38 0.46 0.48 0.50 0.51 0.52
Van-60 60M 0.36 0.42 0.44 0.45 0.45 0.46
Table S5: Top-$K$ accuracy on route test set-n5 (10 000 routes).
DMS SM SM
Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
Van-19 19M 0.39 0.47 0.50 0.52 0.53 0.56
MoE-1/3 38M 0.36 0.42 0.45 0.46 0.46 0.48
Van-60 60M 0.38 0.46 0.48 0.50 0.51 0.52
Van-60 60M 0.36 0.42 0.44 0.45 0.45 0.46
Table S6: Top-$K$ accuracy on subseta of 500 routes from test set-n1.
SM SM
Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
Flash 10M 0.402 0.476 0.522 0.550 0.564 0.582
6x4-12x4b 18M 0.446 0.552 0.598 0.616 0.624 0.640
12x3-12x3 19M 0.452 0.570 0.596 0.618 0.626 0.636
12x4-12x4 22M 0.452 0.552 0.588 0.610 0.620 0.642
6x3-6x3 10M 0.292 0.362 0.384 0.388 0.396 0.402
aThe subset was chosen randomly with a seed of 42. Flash results are provided to establish how the performance on these 500 routes relates to the performance on full test set.
bThe model names follow NxF-KxF format, where N is the number of encoder layers, K is the number of decoder layers, and F is the feed-forward multiplier
Table S7: Top-$K$ accuracy on subseta of 500 routes from test set-n5.
SM SM
Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
Flash 10M 0.374 0.442 0.474 0.484 0.498 0.516
6x4-12x4b 18M 0.400 0.464 0.498 0.520 0.530 0.548
12x3-12x3 19M 0.406 0.472 0.496 0.518 0.524 0.538
12x4-12x4 22M 0.370 0.448 0.472 0.498 0.508 0.526
6x3-6x3 10M 0.222 0.264 0.282 0.288 0.290 0.298
aThe subset was chosen randomly with a seed of 42. Flash results are provided to establish how the performance on these 500 routes relates to the performance on full test set.
bThe model names follow NxF-KxF format, where N is the number of encoder layers, K is the number of decoder layers, and F is the feed-forward multiplier
Table S8: Top-1 Accuracy on n1 routes split by route length
Model 2 3 4 5 6 7 8 9 10
Flash 0.538 0.410 0.323 0.273 0.329 0.308 0.286 0.400 0.000
Explorera 0.480 0.346 0.282 0.266 0.305 0.323 0.286 0.500 0.000
Explorer-XLa,b 0.402 0.304 0.271 0.259 0.296 0.277 0.286 0.600 0.000
Flex (Mono) 0.456 0.318 0.226 0.193 0.183 0.154 0.143 0.300 0.000
Flex (Duo) 0.554 0.426 0.348 0.292 0.362 0.354 0.286 0.500 0.000
Flex (Duo)b 0.471 0.360 0.310 0.278 0.319 0.277 0.286 0.400 0.000
Deepb 0.497 0.395 0.338 0.268 0.272 0.262 0.286 0.600 0.000
Wideb 0.511 0.416 0.344 0.308 0.329 0.292 0.286 0.600 0.000
aSteps not provided during inference.
bSMs not provided during inference.
Table S9: Top-10 Accuracy on n1 routes split by route length
Top10 accuracy on routes with length
Model 2 3 4 5 6 7 8 9 10
Flash 0.763 0.600 0.465 0.385 0.408 0.400 0.333 0.500 0.000
Explorera 0.692 0.546 0.417 0.373 0.376 0.431 0.333 0.500 0.000
Explorer-XLa,b 0.608 0.463 0.387 0.363 0.390 0.400 0.333 0.600 0.000
Flex (Mono) 0.650 0.474 0.325 0.254 0.286 0.231 0.143 0.300 0.000
Flex (Duo) 0.765 0.610 0.469 0.373 0.408 0.385 0.333 0.600 0.000
Flex (Duo)b 0.594 0.460 0.388 0.336 0.371 0.338 0.333 0.600 0.000
Deepb 0.641 0.519 0.442 0.364 0.362 0.323 0.333 0.600 0.000
Wideb 0.647 0.531 0.434 0.378 0.390 0.354 0.333 0.600 0.000
aSteps not provided during inference.
bSMs not provided during inference.
Table S10: Top-1 Accuracy on n5 routes split by route length
Top1 accuracy on routes with length
Model 2 3 4 5 6 7 8 9 10
Flash 0.502 0.401 0.337 0.320 0.331 0.328 0.429 0.692 0.333
Explorera 0.458 0.350 0.300 0.289 0.343 0.288 0.286 0.462 0.333
Explorer-XLa,b 0.395 0.317 0.284 0.303 0.313 0.248 0.405 0.692 0.333
Flex (Mono) 0.436 0.324 0.258 0.211 0.191 0.184 0.190 0.154 0.333
Flex (Duo) 0.549 0.416 0.366 0.336 0.352 0.328 0.476 0.615 0.333
Flex (Duo)b 0.465 0.358 0.324 0.312 0.326 0.344 0.429 0.615 0.333
Deepb 0.474 0.384 0.343 0.313 0.338 0.272 0.381 0.385 0.333
Wideb 0.477 0.399 0.352 0.337 0.366 0.304 0.381 0.615 0.333
aSteps not provided during inference.
bSMs not provided during inference.
Table S11: Top-10 Accuracy on n5 routes split by route length
Top10 accuracy on routes with length
Model 2 3 4 5 6 7 8 9 10
Flash 0.712 0.583 0.479 0.431 0.425 0.400 0.429 0.692 0.333
Explorera 0.654 0.529 0.442 0.404 0.434 0.376 0.405 0.692 0.333
Explorer-XLa,b 0.586 0.469 0.416 0.404 0.423 0.384 0.47 0.692 0.333
Flex (Mono) 0.621 0.461 0.357 0.285 0.297 0.272 0.333 0.462 0.333
Flex (Duo) 0.721 0.594 0.489 0.436 0.444 0.416 0.500 0.692 0.333
Flex (Duo)b 0.581 0.462 0.407 0.383 0.421 0.400 0.452 0.692 0.333
Deepb 0.604 0.507 0.451 0.414 0.430 0.368 0.429 0.615 0.333
Wideb 0.616 0.524 0.451 0.420 0.432 0.328 0.476 0.692 0.333
aSteps not provided during inference.
bSMs not provided during inference.